# Training With “Paraphrasing the Original Text” Teaches LLM to Better Retrieve in Long-context Tasks

### v9

@@ Introduction
Long-context tasks can be divided into two categories: short-dependency and long-dependency (Li et al. 2023). The former means only a small part of the long context is truly needed for the task, while the latter means most parts are needed. Because short-dependency tasks are usually more common and easier to study, for simplicity, in this paper we start with the short-dependency tasks and mainly focus on multi-document-QA, which is one of the most representative long-context tasks and also easy to be constructed.
@@
Intuitively, using CoT (Wei et al. 2023) can help solving composite tasks. Yet, despite using CoT-like prompt, we find many LLMs still perform poorly, which generate wrong answers just in the first step, retrieval, causing the second step to also go wrong, even though some of them can perfectly pass “Needle in a Haystack” test, which represents qualified retrieval ability. While in the other hand, if they retrieve correctly, then the follow-up task becomes easier and the final answer is usually right.
This is either because these models inherently have weak retrieval capabilities, or because their retrieval abilities cannot be fully activated by just designing prompts, which hinders them from accurately locating key information in the long context. This result prompts us to seek a training-based method to enhance the accuracy of models in retrieval-based composite tasks, through teaching them to better retrieve.
Thus, we propose a method based on fine-tuning with specially designed samples to enhance and activate LLMs’ retrieval capability over long contexts. Specifically, for the goal of explicitly separating and highlighting the “retrieval” step, when designing answers of a long-context QA sample, different from normal ways only designing a brief answer, we incorporate an additional part named “original text paraphrasing”, which paraphrases or repeats the original text of the sentences in the context containing relevant information required for answering, as shown in Figure 1. This part corresponds to a direct “retrieval” operation, while maintaining the answers’ quality and coherence, which aims to not only enhance the model’s retrieval capability but also teach the model to use its inherent retrieval capability more effectively.
@@
However, as for composite tasks such as Natural Questions Multi-Document-QA (Liu et al. 2023) and passage retrieval tasks in Longbench (Bai et al. 2023b), it performs much worse. Multi-Document-QA requires model to answer the question based on the context consisting of 20 documents with only 1 document containing truly useful information (i.e. the gold document), and passage retrieval task requires models to answer the corresponding document number given the abstract of one of 30 documents. They can both be considered as composite tasks that require both document understanding and retrieval. As shown in Table 5, it achieves an accuracy of 82% in multi-doc-QA when the context only consists of the gold document, which proves its ability of understanding the document and question is fine. But when the context consists of more documents with the gold document placed in the middle, the accuracy decreases rapidly.
@@
We leverage the power of GPT-4 (OpenAI 2023) to generate high-quality in-context question-answer pairs according to our requirements. Figure 3 shows how we build multi-doc-QA samples with “original text paraphrasing”. For English dataset, we randomly select about 100k samples from the CommonCrawl dataset (Patel 2020) as reference documents for constructing the context, and for Chinese, Wudao 200GB open-source Chinese corpus (BAAI 2023) is used. Then, we let every 100 reference documents form a group, and take the first document from each group as the gold document, based on which, GPT-4 (OpenAI 2023) designs a question-answer pair under the prompt that the relevant original text must be provided in the answer, as follows:
@@
Qwen1.5-4b-Chat SFT w/o ours 65.28 71.86 12.75 19.84 56.5 45.25
@@
Qwen2-7b-Instruct SFT w/ Ziya 65.42 82.4 61.50 19.88 62.50 58.34
@@ Appendix A Appendix
Here are 30 paragraphs from Wikipedia, along with an abstract. Please determine which paragraph the abstract is from. {context} The following is an abstract. {input} Please enter the number of the paragraph that the abstract is from. The answer format must be like P̈aragraph 1,̈ P̈aragraph 2,̈ etc. The answer is:

### v10

@@ Introduction
Long-context tasks can be divided into two categories: short-dependency and long-dependency (Li et al. 2023). The former means only a small part of the long context is truly needed for the task, while the latter means most parts are needed. Because short-dependency tasks are usually more common and easier to study, for simplicity, in this paper, we start with the short-dependency tasks and mainly focus on multi-document-QA, which is one of the most representative long-context tasks and also easy to be constructed.
@@
In this case, it is natural to think that using CoT (Wei et al. 2023) prompt could help. Yet, in our experiments, despite using CoT-like prompt, we find many LLMs still perform poorly, which generate wrong answers just in the first step, retrieval, even though some of them can perfectly pass “Needle in a Haystack” test, which represents qualified retrieval ability. While in the other hand, if they retrieve correctly, then the follow-up task becomes easier and the final answer is usually right.
We posit that this is either because these models inherently have weak retrieval capabilities, or because their retrieval abilities cannot be fully activated by just designing prompts, which hinders them from accurately locating key information in the long context. This prompts us to seek a more effective method to enhance the accuracy of models in retrieval-based composite tasks, through teaching them to better retrieve.
Thus, we propose a method based on fine-tuning with specially designed samples to enhance and activate LLMs’ retrieval capability over long contexts. Specifically, for the goal of explicitly separating and highlighting the “retrieval” step, when designing answers of a long-context QA sample, different from normal ways only designing a brief answer, we incorporate an additional part named “original text paraphrasing”, which paraphrases or repeats the original text of the sentences in the context containing relevant information required for answering, as shown in Figure 1. This part corresponds to a direct “retrieval” operation, while maintaining the answers’ quality and coherence, which aims to not only enhance the model’s retrieval capability but also teach the model to use its inherent retrieval capability more actively.
@@
However, as for composite tasks such as Natural Questions Multi-Document-QA (Liu et al. 2023) and passage retrieval tasks in Longbench (Bai et al. 2023b), it performs much worse. Multi-Document-QA requires model to answer the question based on the context consisting of 20 documents with only 1 document containing truly useful information (i.e. the gold document), and passage retrieval task requires models to answer the corresponding document number given the abstract of one of 30 documents. They can both be considered as composite tasks that require both document understanding and retrieval. As shown in Figure 5, it achieves an accuracy of 82% in multi-doc-QA when the context only consists of the gold document, which proves its ability of understanding the document and question is fine. But when the context consists of more documents with the gold document placed in the middle, the accuracy decreases rapidly.
@@
We leverage the power of GPT-4 (gpt-4-0613) (OpenAI 2023) to generate high-quality in-context question-answer pairs according to our requirements. Figure 3 shows how we build multi-doc-QA samples with “original text paraphrasing”. For English dataset, we randomly select about 100k samples from the CommonCrawl dataset (Patel 2020) as reference documents for constructing the context, and for Chinese, Wudao 200GB open-source Chinese corpus (BAAI 2023) is used. Then, we let every 100 reference documents form a group, and take the first document from each group as the gold document, based on which, GPT-4 (OpenAI 2023) designs a question-answer pair under the prompt that the relevant original text must be provided in the answer, as follows:
@@
In fact, we originally used gpt-3.5-turbo, but when transferred to gpt-4, we found the overall performance was not improved much. Therefore, although intuitively using a more advanced LLM should be better, our method does not necessitate a very strong LLM or a specific version, as long as it can follows the format requirement in most cases.
@@
Qwen1.5-4b-Chat SFT w/o ours 65.28 71.86 12.75 19.84 56.50 45.25
@@
Qwen2-7b-Instruct SFT w/ Ziya 65.42 82.40 61.50 19.88 62.50 58.34
@@
Acknowledgments
This work was supported by National Natural Science Foundation of China (U2336208,82090053,61862002).
We are grateful to Huiqiang Jiang from Microsoft Research Asia for his guidance in writing and revising this paper.
@@ Appendix A Appendix
Here are 30 paragraphs from Wikipedia, along with an abstract. Please determine which paragraph the abstract is from. {context} The following is an abstract. {input} Please enter the number of the paragraph that the abstract is from. The answer format must be like “Paragraph 1”, “Paragraph 2”, etc. The answer is:
@@
Time Consumption
Although our method needs to generate more text, in practice, our method does not increase the time cost too much in long-context scenarios. As shown in Table 7, we compare the inference time cost of our method and the baseline, which indicates there is only a minor increase on time consumption when the context is long.
prompt length baseline ours
3k 1.4 2.5
24k 12.4 15.1
Table 7: The average time consumption (seconds) per QA sample, in different input lengths, in the Multi-Doc-QA task.
Why Previous Training Sample Design is not Optimal: A Conjecture
Ziya-Reader (He et al. 2023) proposed question repetition and index prediction tasks. For the question repetition task, since the fixed position of the question is either at the beginning or end of the context, it is not a segment that needs to be retrieved, hence the task has no contribution to retrieval relevance. In the index prediction task, although the index of the document is directly related to key segments within the context, and thus highly retrieval-relevant, its contribution to the loss in training samples is minimal due to its very short length, which hardly affects overall improvements.